OpenStudy (anonymous):

Perform the indicated operation. (14a^2/10b^2)x(21a^2/15b^2) A.0 B.1 C.b/a

jigglypuff314 (jigglypuff314):

if you mean \[\frac{ 14a ^{2} }{ 10b ^{2} }\times \frac{ 15b ^{2} }{ 21a ^{2} }\] then you can cancel/cross out the variables then simplify to the answer of B. 1